# High-resolution radio imaging of TGSS J1530++1049, a radio galaxy in a dense environment at $z=4$

### v1

@@ 4.2 The nature of the radio features
The fainter feature in the EVN image, component B (Fig. 1) cannot be seen in the e-MERLIN images, neither at $1.5$ GHz (Fig. 2) nor at $5$ GHz (Fig. 3). Its non-detection at $5$ GHz could naturally be explained by the component having a steep spectrum, $\alpha\lesssim-2.4$, rendering it too faint for the sensitivity of the $5$-GHz e-MERLIN observation. However, at the lower frequency, based on its brightness in the higher-resolution EVN image, it should have been visible at a similar significance level as component S. Nevertheless, the rms noise level of the image could have been estimated too optimistically. The $1\sigma$ rms noise level of the $1.5$-GHz e-MERLIN image at the approximate position of component B is $\sim 0.13$ mJy beam-1, about 3 times larger than measured across the whole image, while the largest noise peak is $\sim 0.22$ mJy beam-1. Thus it could be that component B was just below the sensitivity limit of the $1.5$-GHz e-MERLIN observation. Additionally, component B could also be variable. A flux density drop of $\sim 50\,\%$ within $\sim 1.2$ months time in the rest frame of the source ($6$ months elapsed between the EVN and e-MERLIN observations) is able to account for the non-detection of component B with the e-MERLIN at $1.5$ GHz.
@@
The high brightness temperatures of the EVN-detected features indicate AGN origin of the radio emission. While the EVN observation was conducted at a single frequency only, one can use the dual-frequency e-MERLIN measurements to infer the power-law radio spectral index of component A (Fig. 4). If we assume that the $5$-GHz flux density of component N is contained within a compact mas-scale feature, detectable with VLBI, it would imply a spectral index of $\alpha_{\textrm{A}}\approx-1.9$, thus steep radio spectrum. Such spectral index is often related to the lobe-like features seen in radio galaxies, while the cores of radio-loud AGN have usually flat or inverted radio spectra at cm wavelengths (e.g., hovatta_spectralshape, and references therein).
@@
At the position of feature S, there is no sign of any EVN detection, most probably indicating the resolved nature of this component. Moreover, feature S also has a steep radio spectrum indicated by the e-MERLIN data (Fig. 4). Due to the faintness of this feature at $5$ GHz, also indicated by the large relative error of its spectral index, it may have even steeper radio spectrum than the one formally obtained, $\alpha_{\mathrm{S}}\sim-1.7$. Thus, neither of the firmly detected radio features can be regarded as the central radio emitting region of the AGN that is expected to have flat or inverted synchrotron self-absorbed spectrum. The sizes and spectral indices of N and S are more compatible with extended jet-related emissions containing mas-scale compact hot spot such as feature A in the EVN observation.
@@ 4.3 Radio emission at larger scales
According to the $1.4$-GHz NRAO VLA Sky Survey (nvss), J1530++1049 has a flux density of $(7.4\pm 0.5)$ mJy. At the same frequency but with higher resolution, the Faint Images of the Radio Sky at Twenty-centimeters (FIRST) survey measured $(7.25\pm 0.15)$ mJy (first_helfand). Saxena_sample reported a similar flux density value, $(7.5\pm 0.1)$ mJy, measured in a higher-resolution VLA A-configuration observation. More recently, J1530++1049 was detected in the the Rapid Australian Square Kilometre Array Pathfinder (ASKAP) Continuum Survey (RACS with flux densities of $(6.4\pm 0.6)$ mJy, and $(5.6\pm 0.8)$ mJy at $1.37$ GHz, and $1.66$ GHz, respectively (racs-mid; racs-high). All these values exceed the sum of flux densities of components N and S detected with e-MERLIN at close frequencies. However, no self-calibration was performed on the e-MERLIN data, therefore coherence loss can result in lower measured flux density values.
@@
J1530++1049 was detected at $3$ GHz in the three epochs of the VLA Sky Survey (VLASS, vlass_lacy). We downloaded the three images from the Canadian Astronomy Data Centre (CADC555www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/vlass/, accessed on 12 Sep. 2025), and used AIPS to fit them with a single Gaussian. The flux densities agree within the error, their average value is $(2.1\pm 0.5)$ mJy. This exceeds the flux density expected from the spectral indices derived from the e-MERLIN observations (Fig. 4), $\sim 1$ mJy. While a frequency-independent coherence loss of the e-MERLIN data will not alter the obtained spectral index, it can underestimate the predicted $3$-GHz flux density.
@@
Below $1$ GHz, J1530++1049 is brighter, as indicated by the lowest available frequency measurement of the RACS, $888$ MHz (rapid_askap), $(17.1\pm 0.71)$ mJy, and by the TIFR Giant Meterwave Radio Telescope Sky Survey (TGSS, TGSS) at $150$ MHz, $(170\pm 19)$ mJy (Fig. 4). These low-resolution flux density values imply a spectral index of $\alpha_{0.15\mathrm{\,GHz}}^{3\mathrm{\,GHz}}=-1.5\pm 0.1$, but also indicate a flattening of the radio spectrum between $888$ MHz and $150$ MHz, corresponding to $\sim 5$ GHz and $750$ MHz, respectively, in the rest frame of the source at $z=4$.
@@
The $1.4$-GHz radio power of J1530++1049 is between $\sim(3-4)\times 10^{27}\mathrm{\,W\,Hz}^{-1}$, depending on whether the highest $1.4$-GHz flux density and the broad-band spectral index values are used, or whether the flux density and spectral index are taken from the e-MERLIN observations. The linear size of the source is $\sim 5.5$ kpc. It is an order of magnitude fainter and smaller version of the radio galaxy J1420++1205 located at similar redshift ($z=4.026$) and also showing several steep-spectrum features at $100$-mas scale (J1420). The properties of J1530++1049 are similar to those of the high-power compact steep spectrum sources, or medium-sized symmetric objects as shown by e.g. anbaan.
@@ 5 Summary
The northern one could be detected as part of a larger component in our e-MERLIN observation at $1.5$ GHz and $5$ GHz. At larger scales, we detected another steep-spectrum, extended component at $\sim 5.5$ kpc separation from the brightest feature in both e-MERLIN observations. According to the flux densities of the radio features, all have steep power-law spectra. Thus neither of them can be identified as the centre of the AGN. All these features are most probably related to the lobes or hot spots in the AGN, possibly also indicating places of interaction with the surrounding medium.
@@ Acknowledgements.
The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code(s): RSG11. The e-VLBI research infrastructure in Europe was supported by the European Union’s Seventh Framework Programme (FP7/2007-2013) under grant agreement number RI-261525 NEXPReS. The research leading to these results has received funding from the European Commission Horizon 2020 Research and Innovation Programme under grant agreement No. 730562 (RadioNet). e-MERLIN is a National Facility operated by the University of Manchester at Jodrell Bank Observatory on behalf of STFC. Scientific results from data presented in this publication are derived from the e-MERLIN project CY8205. On behalf of the “Interferometric studies of radio-loud active galactic nuclei” project, we are grateful for the possibility to use the HUN-REN Cloud (see Heder2022, https://science-cloud.hu/) which helped us achieve the results published in this paper. This project was supported by the HUN-REN Hungarian Research Network. KÉG and SF received funding from the Hungarian National Research, Development and Innovation Office (NKFIH excellence grant TKP2021-NKTA-64). MVM research has been funded by grant Nr. PID2021-124665NB-I00 by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe”.

### v2

@@ 4.2 The nature of the radio features
The high brightness temperatures of the EVN-detected features indicate the AGN origin of the radio emission. As shown in the previous section, the brighter EVN-detected feature, A, is co-spatial with the e-MERLIN-detected component N. The flux density measured at $1.7$  GHz with the e-MERLIN, $\sim 3.2$ mJy, is slightly higher than the EVN value. This in principle can arise from variability or a resolution issue. The largest recoverable size, determined by the shortest baseline of the array (see Fig. 1), is $\sim 120$ mas for the EVN observation, smaller than the fitted size of component N in the e-MERLIN image ($\sim 152$ mas). Thus, the lower flux density of component A is most likely caused by EVN resolving out some of the flux density contained in the extended feature and thus recovered by e-MERLIN. This indicates that the northern emitting region (component N) is extended but contains a mas-scale compact feature (component A).
@@
While the EVN observation was conducted at a single frequency only, one can use the dual-frequency e-MERLIN measurements to infer a limit on the power-law radio spectral index of component A (Fig. 5). If we assume that all of the $5$-GHz flux density of component N is contained within a compact mas-scale feature, detectable with EVN at $5$ GHz, it would imply a spectral index of $\alpha_{\textrm{A}}\approx-1.9$. This value can be regarded as an upper limit on the spectral index of the compact feature, since similarly to the lower frequency, some amount of the flux density may be resolved out. Such spectral index is often related to the lobe-like features seen in radio galaxies, while the cores of radio-loud AGN have usually flat or inverted radio spectra at cm wavelengths (e.g., Hovatta et al. 2014, and references therein).
The fainter feature in the EVN image, component B (Fig. 2) cannot be seen in the e-MERLIN images, neither at $1.5$ GHz (Fig. 3) nor at $5$ GHz (Fig. 4). Its non-detection at $5$ GHz could naturally be explained by the component having a steep spectrum, $\alpha\lesssim-2.4$, rendering it too faint for the sensitivity of the $5$-GHz e-MERLIN observation. Such extremely steep radio spectra are seen in AGN remnants, old plasma no longer fed by an active jet. Alternatively, old plasma can be re-energized by turbulence or shocks in merging systems (radio phoenices), and show steeply rising radio spectrum towards lower frequencies (e.g., Shulevski et al. 2024, and references therein). However, its compact, few kpc-scale size is not compatible with the known radio phoenices extending for $\sim 100$ kpc size (e.g., Bruno et al. 2025). On the other hand, the estimation of the spectral index, its extreme negative value may be the result of surface brightness sensitivity and angular resolution effects.
Based on its brightness in the higher-resolution EVN image, component B should have been visible at a similar significance level as component S in the $1.5$-GHz e-MERLIN image. Nevertheless, the rms noise level of the image could have been estimated too optimistically. The $1\sigma$ rms noise level of the $1.5$-GHz e-MERLIN image at the approximate position of component B is $\sim 0.13$ mJy beam-1, about $3$ times larger than measured across the whole image, while the largest noise peak is $\sim 0.22$ mJy beam-1. Thus, it could be that component B was just below the sensitivity limit of the $1.5$-GHz e-MERLIN observation. Additionally, component B could also be variable. A flux density drop of $\sim 50\,\%$ within $\sim 1.2$ months time in the rest frame of the source ($6$ months elapsed between the EVN and e-MERLIN observations) is able to account for the non-detection of component B with the e-MERLIN at $1.5$ GHz. If true, such variability is more characteristic for the radio core of the AGN. However, the radio core is not expected to have steep radio spectrum.
@@
At the position of feature S, there is no sign of any EVN detection, most probably indicating the resolved nature of this component. Moreover, feature S also has a steep radio spectrum indicated by the e-MERLIN data (Fig. 5). Due to the faintness of this feature at $5$ GHz, also indicated by the large relative error of its spectral index, it may have even steeper radio spectrum than the one formally obtained, $\alpha_{\mathrm{S}}\sim-1.7$.
Thus, neither of the detected radio features can be clearly identified with the central core of the radio-emitting AGN that is expected to have flat or inverted synchrotron self-absorbed spectrum. The sizes and spectral indices of N and S are more compatible with extended jet-related emissions containing mas-scale compact hot spot such as feature A in the EVN observation. Due to the scarcity of radio data, no firm conclusion can be drawn concerning the nature of component B.
4.3 The radio components and the JWST-detected features
@@
The JWST data furthermore show evidence for perturbed gas, especially close to the location of the southern hotspot. Besides these jet–gas interactions, other sources likely contributing to the strong alignment observed are the ionization by a central, obscured active nucleus and an infalling low-mass starburst galaxy at the southern edge of the field (object C6 in Saxena et al. 2026).
@@ 4.4 Radio emission at larger scales
According to the $1.4$-GHz NRAO VLA Sky Survey (Condon et al. 1998), J1530++1049 has a flux density of $(7.4\pm 0.5)$ mJy. At the same frequency but with higher resolution, the Faint Images of the Radio Sky at Twenty-centimeters (FIRST) survey measured $(7.25\pm 0.15)$ mJy (Helfand et al. 2015). Saxena et al. (2018a) reported a similar flux density value, $(7.5\pm 0.1)$ mJy, measured in a higher-resolution VLA A-configuration observation. More recently, J1530++1049 was detected in the the Rapid Australian Square Kilometre Array Pathfinder (ASKAP) Continuum Survey (RACS with flux densities of $(6.4\pm 0.6)$ mJy, and $(5.6\pm 0.8)$ mJy at $1.37$ GHz, and $1.66$ GHz, respectively (Duchesne et al. 2024, 2025). All these values exceed the sum of flux densities of components N and S detected with e-MERLIN at close frequencies. However, no self-calibration was performed on the e-MERLIN data, therefore coherence loss can result in lower measured flux density values.
@@
J1530++1049 was detected at $3$ GHz in the three epochs of the VLA Sky Survey (VLASS, Lacy et al. 2020). We downloaded the three images from the Canadian Astronomy Data Centre (CADC555www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/vlass/, accessed on 12 Sep. 2025), and used AIPS to fit them with a single Gaussian. The flux densities agree within the error, their average value is $(2.1\pm 0.5)$ mJy. This exceeds the flux density expected from the spectral indices derived from the e-MERLIN observations (Fig. 5), $\sim 1$ mJy. While a frequency-independent coherence loss of the e-MERLIN data will not alter the obtained spectral index, it can underestimate the predicted $3$-GHz flux density.
@@
Below $1$ GHz, J1530++1049 is brighter, as indicated by the lowest available frequency measurement of the RACS, $888$ MHz (Hale et al. 2021), $(17.1\pm 0.71)$ mJy, and by the TIFR Giant Meterwave Radio Telescope Sky Survey (TGSS, Intema et al. 2017) at $150$ MHz, $(170\pm 19)$ mJy (Fig. 5). These low-resolution flux density values imply a spectral index of $\alpha_{0.15\mathrm{\,GHz}}^{3\mathrm{\,GHz}}=-1.5\pm 0.1$, but also indicate a flattening of the radio spectrum between $888$ MHz and $150$ MHz, corresponding to $\sim 5$ GHz and $750$ MHz, respectively, in the rest frame of the source at $z=4$.
@@
The $1.4$-GHz radio power of J1530++1049 is between $\sim(3-4)\times 10^{27}\mathrm{\,W\,Hz}^{-1}$, depending on whether the highest $1.4$-GHz flux density and the broad-band spectral index values are used, or whether the flux density and spectral index are taken from the e-MERLIN observations. The linear size of the source is $\sim 5.5$ kpc. It is an order of magnitude fainter and smaller version of the radio galaxy J1420++1205 located at similar redshift ($z=4.026$) and also showing several steep-spectrum features at $100$-mas scale (Gabányi et al. 2021).
The linear size and radio power of J1530++1049 are similar to those of the medium-sized symmetric objects (MSOs) studied by Fanti et al. (2001) out to a redshift of $\sim 4$. In Fig. 3 in Orienti et al. (2025) [^6] The unit of the horizontal axis is erroneously given in pc, it is in kpc., J1530++1049 could be placed on the evolutionary track shown for objects with jet power of $10^{46}$ erg s-1. MSOs are regarded as young radio sources eventually evolving into large-scale radio galaxies (e.g., Kunert-Bajraszewska et al. 2010; An and Baan 2012; Orienti et al. 2025). Alternatively, their jets can be frustrated, confined within their host galaxies, unable to break through the environment and grow into larger structures. According to Orienti et al. (2025), the progenitor objects of the high-power MSOs are less likely to be influenced by jet instabilities, and expected to eventually evolve into Fanaroff–Riley type II radio galaxies.
Recently, Stanghellini et al. (2025) studied the high-resolution radio morphology and the optical structure of the host galaxies in radio sources containing bright, sub-galactic sized radio jets. They distinguished two groups, those with aligned radio structures likely evolving into radio galaxies in the future and the ones with more complex radio morphologies. They suggested the latter objects reside in merging systems, where the jet interaction with the disturbed interstellar medium, or orbital motion of the jet-emitter AGN in a binary system give rise to rapidly re-orienting jets. There is no clear indication of large bending structures in the radio morphology of J1530++1049. Nevertheless, the three radio-emitting regions, S, B, N (with the hotspot A) are not positioned along a straight line. Additionally, the JWST data revealed a quite disturbed environment around the radio structure, showing the dense region of merging galaxies. Thus, the environment of J1530++1049 may influence its radio structure similarly as found by Stanghellini et al. (2025). More sensitive, high-resolution radio observations conducted at lower frequencies, allowing for the detection of the steep-spectrum jet-related emission, can reveal possible jet disturbances.
@@ 5 Summary
The northern one could be detected as part of a larger component in our e-MERLIN observation at $1.5$ GHz and $5$ GHz. At larger scales, we detected another steep-spectrum, extended component at $\sim 5.5$ kpc separation from the brightest feature in both e-MERLIN observations. According to the flux densities of the radio features, all have steep power-law spectra. Thus neither of them can be identified as the centre of the AGN. All these features are most probably related to the lobes or hot spots in the AGN, possibly also indicating places of interaction with the surrounding medium. Among the detected radio features, the faintest compact component (B), only detected at the $1.7$-GHz EVN observation, is located close to the optical continuum source identified with the AGN. It might be related to the AGN core, however, more sensitive observations are needed to ascertain its nature.
@@ Acknowledgements.
We thank the anonymous referee for their comments which helped us improve our paper. The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code(s): RSG11. The e-VLBI research infrastructure in Europe was supported by the European Union’s Seventh Framework Programme (FP7/2007-2013) under grant agreement number RI-261525 NEXPReS. The research leading to these results has received funding from the European Commission Horizon 2020 Research and Innovation Programme under grant agreement No. 730562 (RadioNet). e-MERLIN is a National Facility operated by the University of Manchester at Jodrell Bank Observatory on behalf of STFC. Scientific results from data presented in this publication are derived from the e-MERLIN project CY8205. On behalf of the “Interferometric studies of radio-loud active galactic nuclei” project, we are grateful for the possibility to use the HUN-REN Cloud (see Heder2022, https://science-cloud.hu/) which helped us achieve the results published in this paper. This project was supported by the HUN-REN Hungarian Research Network. KÉG and SF received funding from the Hungarian National Research, Development and Innovation Office (NKFIH excellence grant TKP2021-NKTA-64). MVM research has been funded by grant Nr. PID2021-124665NB-I00 by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe”.